A sail boat approaches a dock at.500 m/s. A piece
Answer #1

Using

:ycxS

We get R = 19.76 m

time of flight =

= 2.756s

We can calculate maximum height reached by projectile above 9m scene using

h = v2sin2\theta/2g = 5.35 m

So maximum height above ground = 9+5.35 = 14.35m

Speed along vertical = vinitial -gt = 12.5sin 55 - 9.8*2.756 = -16.77m/s

Speed along horizontal = vcos\theta = 12.5cos55 = 7.17 m/s

so vnet = 18.24m/s

For angle \theta =tan-1(16.77/7.17) = 66.85o below horizontal